Could someone please help me with this equation.
Find two consecutive positive integers such that the sum of their squares is 85.

Answer by edjones(8007)

let x, x+1, be the integers
x^2+(x+1)^2=85
x^2+x^2+2x+1=85
2x^2+2x-84=0 subtract 85 from each side and combine terms.
2(x^2+x-42)=0 factor
2(x+7)(x-6)=0
x=-7 and x=6
so x=6 because the problem calls for positive integers.
Check:
6^2+7^2=36+49=85
Ed